# Learning in Deep Factor Graphs with Gaussian Belief Propagation

### v1

@@ Appendix A Factor to Variable Message Update Optimisation
Appendix B Toy Experiment Details
B.1 XOR
@@
We run GBP in the training graph (with $4$ input/output observations) for $600$ iterations. We then fix the parameters, remove the softmax class observation layer and run GBP for $300$ iterations on a grid of $20\times 20$ test input points. The last layer activation variables are then treated as the logits for class predictions. At both train and test time, we use a damping factor of $0.8$ and a dropout of $0.5$ in the factor to variable message updates from the dense factors.
@@
Inc. bias N/A
Weight prior $\sigma$ $3.0$ $3.0$ N/A
Activation prior $\sigma$ $5.0$ $2.0$ N/A
Input obs. $\sigma$ $0.02$ N/A N/A
Class obs. $\sigma$ N/A N/A $0.1$
Dense recon. $\sigma$ $0.1$ $0.1$ N/A
Activation function $g(\cdot)$ Leaky ReLU Linear N/A
@@ B.2 Regression
$z_{i}$ $=5\cdot\sin\left(6.7\cdot x_{i}\right)+\left(10\cdot x\right)^{2}*0.15+\epsilon_{i}~{},$ (24)
@@
To fit the data, we use a similar MLP-inspired factor graph architecture to that illustrated in Fig. 1(a). However, this architecture has only $1$ input variable and $1$ output variable. We use $16$ units in the hidden layer and a Leaky ReLU activation in the first-layer dense factor. The full model is described in Table 2.
We run GBP in the training graph (with $90$ input/output observations) for $2000$ iterations. We then fix the parameters, remove the output observation layer and run GBP for $1000$ iterations on $225$ uniformly spaced test input points. The last layer activation variables are then treated as the predictions for these query points. At both train and test time we use a damping factor of $0.8$ and a dropout of $0.6$ in the factor to variable message updates from the dense factors.
@@
Inc. bias N/A
Weight prior $\sigma$ $7.0$ $7.0$ N/A
Activation prior $\sigma$ $5.0$ $1.0$ N/A
Input obs. $\sigma$ $0.02$ N/A N/A
Output obs. $\sigma$ N/A N/A $0.03$
Dense recon. $\sigma$ $7\times 10^{-3}$ $7\times 10^{-3}$ N/A
Activation function $g(\cdot)$ Leaky ReLU Linear N/A
@@
Appendix C Video Denoising Experiment Details
C.1 Single Layer Convolutional Factor Graph
We use the transposed convolution model described in Table C1 for both per-frame and continual learning experiments. We run for $300$ GBP iterations on each frame with damping factor of $0.8$ and dropout factor of $0.6$ applied to the factor to variable messages. We used robust factor energies similar to the Tukey loss (Tukey:1960): quadratic within a Mahalanobis distance of $N_{\mathrm{rob}}$ from the mean, and flat outside.
@@ C.1 Single Layer Convolutional Factor Graph
Table C1: The single layer convolutional factor graph model used for the video denoising experiments. $N_{\mathrm{rob}}$ is the “robust threshold”: the Mahalanobis distance beyond which the factor energy is flat rather than quadratic.
C.1.1 Preventing overfitting in single-layer model
@@
C.2 Five Layer Convolutional Factor Graph
We use the transpose convolution model described in Table C2 for both per-frame and continual learning experiments. We run for $500$ GBP iterations on each frame with damping factor of $0.8$ and dropout factor of $0.6$ applied to the factor to variable messages. We used robust factor energies similar to the Tukey loss (Tukey:1960): quadratic within a Mahalanobis distance of $N_{\mathrm{rob}}$ from the mean, and flat outside.
To increase the spatial extent of the activations, we use upsampling layers within which each output $x_{l}^{(a,b,c)}$ connects to a $K\times K$ patch of the input $\mathbf{X}_{l-1}^{(a,b,c)}$ via a factor with energy
$E_{\mathrm{upsample}}$ $=\frac{1}{2\sigma^{2}}\sum_{i,j=-\lfloor K/2\rfloor{}}^{\lfloor K/2\rfloor{}}\left(x_{l-1}^{(a+i,b+j,c)}-x_{l}^{(a,b,c)}\right)^{2}~{}.$ (30)
@@ C.2 Five Layer Convolutional Factor Graph
Number of filters 4 N/A 8 N/A 8
Inc. bias N/A N/A
@@
Recon $N_{\mathrm{rob}}$ 2.5 N/A N/A N/A N/A
Weight prior $\sigma$ 0.15 N/A 0.3 N/A 0.3
Activation prior $\sigma$ 0.5 N/A 0.5 N/A 0.5
Pixel obs. $\sigma$ 0.2 N/A N/A N/A N/A
Pixel obs $N_{\mathrm{rob}}$ 0.2 N/A N/A N/A N/A
Activation func. $g(\cdot)$ Linear N/A Leaky ReLU N/A Leaky ReLU
Table C2: The five-layer convolutional factor graph model used for the video denoising experiments. $N_{\mathrm{rob}}$ is the “robust threshold”: the Mahalanobis distance beyond which the factor energy is flat rather than quadratic.
C.3 Pairwise Factor Graph
The factor hyperparameters for the pairwise smoother baseline are given in Table C3. We denoise by running $200$ GBP iterations on each frame with damping factor of $0.7$ applied to the factor to variable messages. We used robust factor energies similar to the Tukey loss (Tukey:1960): quadratic with a Mahalanobis distance of $N_{\mathrm{rob}}$ from the mean, and flat outside.
@@ C.3 Pairwise Factor Graph
Table C3: The pairwise smoothing baseline model used for the video denoising experiments. $N_{\mathrm{rob}}$ is the “robust threshold”: the Mahalanobis distance beyond which the factor energy is flat rather than quadratic.
Appendix D Denoised Video Frame Example
@@
(a) Clean image
(b) Corrupted image
(c) Per-frame learning, single layer
(d) Continual learning, single layer
(e) Per-frame learning, five layer
(f) Continual learning, five layer
(g) Pairwise smoothing
Appendix E MNIST experiment details
E.1 Deep Gaussian Factor Graph
@@
The final convolutional factor graph model is summarised in Table E1. To produce the results shown in Fig. 4, we train with continual learning, with a batchsize of $50$ and run $800$ GBP iterations on each batch. At test time, we fix the parameters and run GBP for $500$ iterations per test batch of $200$ examples. We apply a damping factor of $0.9$ and a dropout factor of $0.5$ to the factor to variable messages at both train and test time. Note that each iteration includes the updating of the messages in each layer, sweeping from image layer to classification head, and then sweeping back to image.
@@
Num. filters 8 N/A N/A N/A
Kernel size $5\times 5$ $2\times 2$ N/A N/A
Dense num. inputs N/A N/A 1152 N/A
Dense num. outputs N/A N/A 10 N/A
Inc. bias N/A N/A
Weight prior $\sigma$ 0.04 N/A 0.3 N/A
Activation prior $\sigma$ 1.0 1.0 2.0 N/A
Pixel obs. $\sigma$ 0.03 N/A N/A N/A
Recon. $\sigma$ 0.02 0.02 0.02 N/A
Activation func. $g(\cdot)$ Leaky ReLU Linear Linear N/A
Class observation $\sigma$ N/A N/A N/A 0.01
@@
Table E1: The convolutional factor graph model used for the MNIST experiment. The dimensions of the layers are chosen such that no padding is necessary. Note that “Recon $\sigma$” denotes the strength of the factors which connect one layer to the next. ${}^{\dagger}$oftmax layer only included at training time, when class observation is available.
E.2 Linear Classifier Baseline
@@
As a baseline we used a linear classifier trained by backpropagation of the cross-entropy loss. We tuned the following on a randomly sampled validation set comprising $9,000$ training set examples:
  • Training:
    • Optimiser type: Adam (kingma2014adam) vs SGD
    • Optimiser step size
    • Number of epochs
  • Regularisation: $L2$ coefficient.
We tuned a separate linear classifier for each training set size. As with the factor graph model, we used a batchsize of $50$ for all experiments in Fig. 4.
E.3 Asynchronous Training
@@
To test the robustness of our method to asynchronous training regimes, we evaluated training the model in Table E1 with random layer schedules. At each iteration, we randomly sample $4$ layers, and update the messages within each of the sampled layers in the sampled order. Note that we sample with replacement, meaning that some layers may not be updated at all during that iteration, and some may be updated multiple times.
As each iteration updates the layers in a different order, and the different layers comprise different shapes and operations, this asynchronous training cannot be readily compiled. As a result, it is much slower to run. To mitigate this, we train with a larger batch size of $200$ (compared to the batch size of $50$ used in the experiments for Fig. 4). We also train the model with synchronous forward/backward sweeps, and a batch size of $200$. The results reported in Section 4.3 correspond to $3$ repeats with different random seeds.
E.4 Dependence on number of iterations
@@
We seek to understand how the performance of our models depends on different levels of compute at train and test time. We train replicas of the convolutional factor graph model (summarised in Table E1) on MNIST, each with a different number of GBP iterations per training batch. We then evaluate the test accuracy of each of the trained models multiple times: for differing numbers of test time GBP iterations. All configurations used a batch size of $200$. We ran our experiments with $50,100,200,400,800,1600$ iterations per batch at both train time and test time. However, we found training with $50$ iterations per batch diverged, so this is not included in the results.
@@
The results are presented in Fig. E1. They show that good test time performance can be achieved with relatively few GBP iterations per batch ($\sim 100$), as long as a sufficient number of iterations is run at test time ($\geqslant 200$).
@@
Figure E1: Dependence of MNIST test accuracy on the number of GBP iterations at train and test time.
auqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ Appendix B Factor to Variable Message Update Optimisation
Appendix C Toy Experiment Details
C.1 XOR
@@
We run GBP in the training graph (with $4$ input/output observations) for $600$ iterations. We then fix the parameters, remove the softmax class observation layer and run GBP for $300$ iterations on a grid of $20\times 20$ test input points. The last layer activation variables are then treated as the logits for class predictions. At both train and test time, we use a damping factor of $0.7$ applied to the factor to variable message updates from the dense factors.
@@
Inc. bias -
Weight prior $\sigma$ $3.0$ $3.0$
Activation prior $\sigma$ $5.0$ $2.0$
Input obs. $\sigma$ $0.02$
Class obs. $\sigma$ $0.1$
Dense recon. $\sigma$ $0.1$ $0.1$
Activation function $g(\cdot)$ Leaky ReLU Linear -
@@ C.2 Regression
$z_{i}$ $=5\cdot\sin\left(6.7\cdot x_{i}\right)+\left(10\cdot x_{i}\right)^{2}\cdot 0.15+\epsilon_{i}~{},$ (24)
@@
To fit the data, we use a similar MLP-inspired factor graph architecture to that illustrated in Fig. 1(a). However, this architecture has only $1$ input variable and $1$ output variable. We use $16$ units in the hidden layer and a sigmoid activation in the first-layer dense factor. The full model is described in Table 3.
We run GBP in the training graph (with $90$ input/output observations) for $4000$ iterations. We then fix the parameters, remove the output observation layer and run GBP for $1000$ iterations on $225$ uniformly spaced test input points. The last layer activation variables are then treated as the predictions for these query points. At both train and test time we use a damping factor of $0.8$ and a dropout of $0.6$ in the factor to variable message updates from the dense factors.
@@
Inc. bias -
Weight prior $\sigma$ $6.0$ $1.5$
Activation prior $\sigma$ $10.0$ $3.0$
Input obs. $\sigma$ $0.02$
Output obs. $\sigma$ $0.05$
Dense recon. $\sigma$ $5\times 10^{-3}$ $1\times 10^{-2}$
Activation function $g(\cdot)$ Sigmoid Linear -
@@
Appendix D Video Denoising Experiment Details
D.1 Single Layer Convolutional Factor Graph
We use the transposed convolution model described in Table D1 for both per-frame and continual learning experiments. We run for $300$ GBP iterations on each frame with damping factor of $0.8$ and dropout factor of $0.6$ applied to the factor to variable messages. We used robust factor energies similar to the Tukey loss (Tukey:1960): quadratic within a Mahalanobis distance of $N_{\mathrm{rob}}$ from the mean, and flat outside.
@@ D.1 Single Layer Convolutional Factor Graph
Table D1: The single layer convolutional factor graph model used for the video denoising experiments. $N_{\mathrm{rob}}$ is the “robust threshold”: the Mahalanobis distance beyond which the factor energy is flat rather than quadratic.
D.1.1 Preventing overfitting in single-layer model
@@
D.2 Five Layer Convolutional Factor Graph
We use the transpose convolution model described in Table D2 for both per-frame and continual learning experiments. We run for $500$ GBP iterations on each frame with damping factor of $0.8$ and dropout factor of $0.6$ applied to the factor to variable messages. We used robust factor energies similar to the Tukey loss (Tukey:1960): quadratic within a Mahalanobis distance of $N_{\mathrm{rob}}$ from the mean, and flat outside.
To increase the spatial extent of the activations, we use upsampling layers within which each output $x_{l}^{(a,b,c)}$ connects to a $K\times K$ patch of the input $\mathbf{X}_{l-1}^{(a,b,c)}$ via a factor with energy (15).
@@ D.2 Five Layer Convolutional Factor Graph
Number of filters 4 - 8 - 8
Inc. bias - N/A
@@
Recon $N_{\mathrm{rob}}$ 2.5 - - - -
Weight prior $\sigma$ 0.15 - 0.3 - 0.3
Activation prior $\sigma$ 0.5 - 0.5 - 0.5
Pixel obs. $\sigma$ 0.2 - - - -
Pixel obs $N_{\mathrm{rob}}$ 0.2 - - - -
Activation func. $g(\cdot)$ Linear - Leaky ReLU - Leaky ReLU
Table D2: The five-layer convolutional factor graph model used for the video denoising experiments. $N_{\mathrm{rob}}$ is the “robust threshold”: the Mahalanobis distance beyond which the factor energy is flat rather than quadratic.
D.3 Pairwise Factor Graph
The factor hyperparameters for the pairwise smoother baseline are given in Table D3. We denoise by running $200$ GBP iterations on each frame with damping factor of $0.7$ applied to the factor to variable messages. We used robust factor energies similar to the Tukey loss (Tukey:1960): quadratic with a Mahalanobis distance of $N_{\mathrm{rob}}$ from the mean, and flat outside.
@@ D.3 Pairwise Factor Graph
Table D3: The pairwise smoothing baseline model used for the video denoising experiments. $N_{\mathrm{rob}}$ is the “robust threshold”: the Mahalanobis distance beyond which the factor energy is flat rather than quadratic.
Appendix E Denoised Video Frame Example
(a) Clean image
(b) Corrupted image
(c) Per-frame learning, single layer
(d) Continual learning, single layer
(e) Per-frame learning, five layer
(f) Continual learning, five layer
(g) Pairwise smoothing
@@ Appendix E Denoised Video Frame Example
Appendix F MNIST experiment details
F.1 Convolutional Factor Graph
@@
The final convolutional factor graph model is summarised in Table D4. Pixel variables are fixed at their observed values. To produce the results shown in Fig. 4, we train with continual learning, with a batchsize of $50$ and run $500$ GBP iterations on each batch. At test time, we fix the parameters and run GBP for $300$ iterations per test batch of $200$ examples. We apply a damping factor of $0.9$ and a dropout factor of $0.5$ to the factor to variable messages at both train and test time. Note that each iteration includes the updating of the messages in each layer, sweeping from image layer to classification head, and then sweeping back to image.
@@
Num. filters 16 - - -
Kernel size $5\times 5$ $2\times 2$
Dense num. inputs - - 2304 -
Dense num. outputs - - 10 -
Inc. bias - -
Weight prior $\sigma$ 0.1 - 0.15 -
Activation prior $\sigma$ 3.0 3.0 2.0 -
Recon. $\sigma$ 0.01 0.01 0.01 -
Activation func. $g(\cdot)$ Leaky ReLU Linear Linear -
Class observation $\sigma$ 0.01
@@
Table D4: The convolutional factor graph model used for the MNIST experiment. The dimensions of the layers are chosen such that no padding is necessary. Note that “Recon $\sigma$” denotes the strength of the factors which connect one layer to the next. ${}^{\dagger}$oftmax layer only included at training time, when class observation is available.
F.2 Linear Classifier Baseline
@@
As a baseline we used a linear classifier trained with Adam to minimise the cross-entropy loss. We tuned the step size and number of epochs on a randomly sampled validation set comprising $9,000$ training set examples. We tuned a separate linear classifier configuration for each training set size. As with the factor graph model, we used a batchsize of $50$ for all experiments in Fig. 4.
F.3 CNN + Replay Buffer Baseline
We train a CNN with Adam as a comparison against our convolutional factor graph trained with GBP Learning. To make this comparison as fair as possible we
  • Use a CNN architecture as close as possible to our factor graph (Table D4). The CNN architecture is summarised in Table D5.
  • Train for only one epoch, but equip the CNN with a FIFO replay buffer to reduce forgetting. A fixed number of elements from each batch are randomly selected and added to the buffer. The size of the buffer necessary to match the performance of GBP Learning then provides an indication as to the efficacy of our continual learning approach. We evaluate buffers of various sizes. To allow comparison across different dataset sizes, we express buffer size as a fraction of the training set size.
We tune the following hyperparameters of the CNN + replay buffer method, finding a different configuration for each combination of buffer size (as fraction of training set size) and training set size:
  • Number of elements in each batch added to the buffer
  • Number of steps to take on each training set batch
  • Number of steps to take on batches sampled from the replay buffer, for each training set batch
  • Step size for training set batch updates
  • Step size for replay buffer updates.
These hyperparameters were selected based on a validation set of $9,000$ examples, sampled uniformly from the training set. The Adam optimiser kingma2014adam was used for all parameter updates.
Layer # 1 2 3
Layer type Conv. Max pool Dense
Num. filters 16 - -
Kernel size $5\times 5$ $2\times 2$
Dense num. inputs - - 2304
Dense num. outputs - - 10
Inc. bias -
Table D5: The baseline CNN architecture.
F.4 Asynchronous Training
@@
To test the robustness of our method to asynchronous training regimes, we evaluated training the model in Table D4 on MNIST using random layer schedules. At each iteration, we randomly sample $4$ layers, and update the messages within each of the sampled layers in the sampled order. Note that we sample with replacement, meaning that some layers may not be updated at all during that iteration, and some may be updated multiple times.
The progression of test accuracy for both asynchronous and synchronous training is presented in Fig. D3. Both regimes exhibit similar performance throughout training, suggesting that GBP Learning can be executed in a distributed and asynchronous manner with little loss in performance.
Figure D3: The evolution in test accuracy over the course of training, for a convolutional factor graph (Table D4) trained on MNIST, with different message schedules: i) synchronous forward/backward sweeps and ii) random layer ordering. Intervals represent ±$\pm 1$ standard error either side of the mean, over $5$ random seeds.
F.5 Dependence on number of iterations
@@
We seek to understand how the performance of our models depends on different levels of compute at train and test time. We train replicas of the convolutional factor graph model (summarised in Table D4) on MNIST, each with a different number of GBP iterations per training batch. We then evaluate the test accuracy of each of the trained models multiple times: for differing numbers of test time GBP iterations. All configurations used a batch size of $50$ at train time and $200$ at test time. We ran our experiments with $50,100,200,400,800,1600$ iterations per batch at both train time and test time.
@@
The results are presented in Fig. D4. They show that good test time performance can be achieved with relatively few GBP iterations per training batch ($\sim 200$), as long as a sufficient number of iterations is run at test time ($\geqslant 200$). However, the best performance is achieved with $1600$ iterations per train batch.
@@
Figure D4: Dependence of MNIST test accuracy on the number of GBP iterations at train and test time. The right-hand plot covers the range of iterations marked by the red box in the left-hand plot, with the colours rescaled accordingly.